# When Distance Distracts: Representation Distance Bias in BT-Loss for Reward Models

### v1

@@ 5 Conclusion
In this study, we analyze the widely used Bradley-Terry loss for reward modeling and identify a key limitation in its update dynamics. We show that update magnitude depends jointly on (1) prediction error, and (2) representation distance between the response pair, with the latter introducing biased learning signals that are misaligned with model performance. To this end, we propose NormBT, as a lightweight modification to BT-loss through pair-wise normalization. Experiments across various base models and training datasets show consistent performance gains of NormBT over the standard BT objective. Our findings provide insights into extracting faithful and efficient learning signals, thereby facilitating preference modeling and LM alignment.
6 Ethics Statement
Our work on reward modeling is committed to responsible AI development and adheres to standard academic and ethical practices. This particular project does not involve human subjects or raise concerns regarding data privacy, bias, or fairness in its current scope. Our research focuses on foundational architectural and training methodologies, with no direct application to the creation of sensitive or harmful queries. We are dedicated to ensuring that our research contributes to the safe and beneficial advancement of AI and are actively exploring methods to detect and prevent malicious applications of LLM alignment.
7 Reproducibility Statement
To ensure reproducibility of our results, we provide the following resources: (1) complete implementation details and hyperparameters are described in Sec 3 and Appendix B; (2) all benchmarks and models used in our experiments are publicly available and properly cited with access information provided in Sec 3; and (3) source code will be made available upon publication to facilitate replication of our experimental results.
@@ Appendix A Limitations & Future Work
Evaluation. Our comparison focuses on BT-based variants. Techniques involving regularization or auxiliary objectives may alter gradient dynamics beyond Eq. 7, and exploring how NormBT interacts with such methods is an interesting direction. We use Best-of-N to mirror downstream RLHF usage; full policy-level evaluations (schulman2017proximalpolicyoptimizationalgorithms; ahmadian2024basicsrevisitingreinforcestyle) would offer a more comprehensive assessment.
@@ Appendix B Implementation Details
  • BT + Margin. (touvron2023llama2openfoundation) For the Unified-Feedback dataset, which provides ratings for the responses, we extract a ground-truth margin $m$ (i.e., difference in ratings) and incorporate it into the BT loss. In this formulation, the margin is introduced inside the log-sigmoid:
    $\mathcal{L}_{\text{Margin}}(\theta)$ $=-\mathbb{E}_{(x,y_{w},y_{l},m)\sim D}\left[\log\,\sigma\big(r_{w}-r_{l}-{\color[rgb]{0,0,1}m}\big)\right].$ (14)
@@
  • BT + Margin (outside)(wang2025helpsteer2preferencecomplementingratingspreferences) In the second formulation, the margin is applied outside of log-sigmoid as a multiplicative weight on the per-sample loss:
    $\mathcal{L}_{\text{Margin(out)}}(\theta)$ $=-\mathbb{E}_{(x,y_{w},y_{l},m)\sim D}\left[{\color[rgb]{0,0,1}m}\cdot\log\,\sigma\big(r_{w}-r_{l}\big)\right].$ (15)
@@
  • BT + Label Smoothing. (liu2024skyworkrewardbagtricksreward) A widely used regularization technique in classification models, where soft labels replace hard binary targets:
    $\mathcal{L}_{\text{LS}}(\theta)$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[{\color[rgb]{0,0,1}(1-\alpha)}\log\,\sigma\big(r_{w}-r_{l}\big)+{\color[rgb]{0,0,1}\alpha\big(r_{l}-r_{w}\big)}\right].$ (16)
@@
    $\nabla_{\theta}\mathcal{L}_{\text{LS}}=\big(\sigma(d)-{\color[rgb]{0,0,1}(1-\alpha)}\big)\cdot\big(\nabla_{\theta}r_{w}-\nabla_{\theta}r_{l}\big),$ (17)
@@
    $=\Bigl(\underbrace{1-\sigma(d)}_{\text{BT magnitude}}-{\color[rgb]{0,0,1}\alpha}\Bigr)\,\bigl\|\,\nabla_{\theta}r_{w}-\nabla_{\theta}r_{l}\,\bigr\|.$ (18)
@@
Training Details.
We implement all methods based on transformers (wolf2020huggingfacestransformersstateoftheartnatural) and trl (vonwerra2022trl), developing from the work of GRM (yang2024regularizing). To use the Unified-Feedback dataset, we downsample the training data from the “all” set. All reward models are trained with the default reward head, as a linear layer with shape (hidden size, 1). For NormBT, the normalization strength $\alpha$ is set to 1 by default. For label smoothing, $\alpha$ is set to 0.1 by default. For all BT baselines, we conduct an extensive grid search over learning rates and report the best-tuned results. For BT-based variants (margin, margin-out, label smoothing), we search over a range of learning rates centered around the best value found for BT and report the best configuration. All models are trained for 1 epoch with full parameter tuning. We truncate the inputs over 4096 tokens, and an effective batch size of 256 with gradient accumulation. Our experiments are conducted using NVIDIA RTX A6000 49G.
@@
Appendix C Intuition: Large- vs. Small-Distance Pairs
Small representation distance can arise from two reasons: (1) uninformative or noisy pairs, such as near-duplicate responses that convey little preference signal; or (2) difficult but meaningful pairs, where the backbone fails to separate two responses despite clear human preference signal (e.g., correct vs. subtly flawed code). As illustrated in Figure 2, such case (2) pairs are common in practice, and Figure 7 provides concrete examples.
In our datasets (UnifiedFeedback and Skywork-80K), most pairs are human-annotated or carefully curated, and following yang2024regularizing, pairs with identical ratings are filtered out. This substantially reduces the prevalence of case (1), making small-distance pairs far more likely to reflect case (2) of the difficult, fine-grained distinctions that RM should learn. Viewed through this lens, NormBT effectively upweights hard, meaningful pairs whose gradients would otherwise be suppressed due to representation limitations in the current model.
Appendix D Active Learning Approaches
The above intuition also clarifies how our analysis relates to active learning approaches such as PILAF (feng2025pilafoptimalhumanpreference) and ARM (shen2025revivingclassicsactivereward), which prioritize large-distance pairs to improve sample efficiency. While this appears at odds with our motivation, the two settings address fundamentally different problems.
Active-learning methods focus on sample selection: given a noisy pool of model-generated responses, identify pairs that are maximally informative. In this setting, many candidates reflect case (1) from Appendix C as uninformative or trivial pairs that provide little preference signal. Moreover, ARM adopts an embedding-based RM with a frozen LLM backbone, meaning $\Phi$ is assumed reliable and representation learning is not part of training. Under these assumptions, difficult but meaningful small-distance pairs cannot be corrected, thus selecting large-distance pairs becomes a principled strategy for sample efficiency.
In contrast, sequence-classifier RMs jointly train the backbone and reward head. The assumption of fixed, reliable $\Phi$ no longer holds, especially as the model newly adapts to the task. Therefore, a small-distance pair could arise because the backbone has not yet learned to separate meaningful distinctions (i.e., case (2)), rather than a lack of meaningful preference information. These are precisely the pairs where the reward model needs stronger updates. Yet under standard BT loss, both case (1) and case (2) receive small gradients, creating a learning bottleneck.
This distinction is amplified in our datasets. UnifiedFeedback and Skywork-80K contain human-annotated or carefully curated pairs, and identical-rating pairs are filtered out following yang2024regularizing. This substantially removes case (1) examples, making small-distance pairs far more likely to represent meaningful but difficult decisions. In this regime, prioritizing large-distance pairs (as active-learning methods do) suppresses the critical fine-grained distinctions.
Our empirical results also reflect this: active-learning style D-opt scaling underperforms in our setting (Table 2), since the difficult, fine-grained distinctions driving performance in Reasoning are further suppressed. In contrast, NormBT improves these regions while maintaining competitive performance elsewhere. However, we also note that our D-opt scaling uses a heuristic continuous relaxation for fair comparison using the same dataset and ordering. This differs from its original use as a sample-selection criterion, and we include it only to highlight the conceptual differences between active-learning approaches and NormBT in the context of reward model optimization.
Appendix E Experiment Error Analysis
To assess the robustness of our main results reported in Table 1, we conduct an expanded evaluation on five random seeds for each method. Using a 40K subset of the Unified-Feedback dataset, we train both NormBT and the BT baseline following the same training configuration described in Appendix B. We compare these 10 models to illustrate the performance improvement of NormBT over the BT baseline. The results reported in Table 4 and Figure 6 are consistent with the main experiments, showing higher average scores across the majority of RewardBench and particularly strong gains in the Reasoning category.
Reward Model Average Chat Chat Hard Safety Reasoning
BT (Baseline) 68.15 94.69 37.48 75.11 65.34
NormBT (Ours) 69.62 95.03 38.00 74.95 70.50
Table 4: Comparison of Ten Models. Mean performance on RewardBench over models trained on five random seeds aligns with the main results reported in Table 1.
@@ Appendix E Experiment Error Analysis
Figure 6: Comparison on Ten Models. Detailed performance of models trained on five random seeds with standard errors on (a) RewardBench categories, and (b) Unified-Feedback Evaluation.
Appendix F Additional Examples
@@
Figure 7: Small-Distance Pairs. Example response pairs from RewardBench, where NormBT predicts correctly on while BT Baseline does not. These pairs fall within the “Small-Distance” range. Response pairs from three different prompts are shown, where left shows the chosen the right shows the rejected response.
@@ Appendix G Representation Scale & EMA
Appendix H Full vs. Proxy Representation Distance
(a) gemma-2b-it
(b) Llama-3.2-3b-Insturct
(c) Llama (with training)

### v2

@@ 5 Conclusion
In this study, we analyze the widely used Bradley-Terry loss for reward modeling and identify a key limitation in its update dynamics. We show that update magnitude depends jointly on (1) prediction error, and (2) representation distance between the response pair, with the latter introducing biased learning signals that are misaligned with model performance. To this end, we propose NormBT, as a lightweight modification to BT-loss through pair-wise normalization, which rescales gradient contributions and ensures updates are driven by prediction errors. Experiments across various base models and training datasets show consistent performance gains of NormBT over the standard BT objective. Our findings provide insights into extracting faithful and efficient learning signals, thereby facilitating preference modeling and LM alignment.
Impact Statement
This work studies the update dynamics of Bradley-Terry (BT) loss for training reward models and proposes NormBT as an effective normalization scheme. By improving learning from fine-grained preference pairs, our method may lead to more reliable reward signals in downstream applications such as RLHF, potentially improving helpfulness and reducing failure modes where subtle reasoning errors are under-penalized. NormBT does not by itself address issues of data quality, annotator bias, or the broader risks associated with deploying high-capability language models. Therefore, we encourage practitioners to pair reward-model improvements with careful dataset curation and bias evaluation when applying the method in real-world systems.
@@ Appendix A Limitations & Future Work
Evaluation. Our comparison focuses on BT-based variants. Techniques involving regularization or auxiliary objectives may alter gradient dynamics beyond Eq. 7, and exploring how NormBT interacts with such methods is an interesting direction. We use Best-of-N to mirror downstream RLHF usage; full policy-level evaluations (Schulman et al., 2017; Ahmadian et al., 2024) would offer a more comprehensive assessment.
@@ Appendix B Implementation Details
  2. 2.
    BT + Margin. (Touvron et al., 2023) For the Unified-Feedback dataset, which provides ratings for the responses, we extract a ground-truth margin $m$ (i.e., difference in ratings) and incorporate it into the BT loss. In this formulation, the margin is introduced inside the log-sigmoid:
    $\mathcal{L}_{\text{Margin}}(\theta)$ $=-\mathbb{E}_{(x,y_{w},y_{l},m)\sim D}\left[\log\,\sigma\big(r_{w}-r_{l}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m}\big)\right].$ (14)
@@
  3. 3.
    BT + Margin (outside)(Wang et al., 2025) In the second formulation, the margin is applied outside of log-sigmoid as a multiplicative weight on the per-sample loss:
    $\mathcal{L}_{\text{Margin(out)}}(\theta)$ $=-\mathbb{E}_{(x,y_{w},y_{l},m)\sim D}\left[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m}\cdot\log\,\sigma\big(r_{w}-r_{l}\big)\right].$ (15)
@@
  4. 4.
    BT + Label Smoothing. (Liu et al., 2024a) A widely used regularization technique in classification models, where soft labels replace hard binary targets:
    $\mathcal{L}_{\text{LS}}(\theta)$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha)}\log\,\sigma\big(r_{w}-r_{l}\big)+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha\big(r_{l}-r_{w}\big)}\right].$ (16)
@@
    $\nabla_{\theta}\mathcal{L}_{\text{LS}}=\big(\sigma(d)-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha)}\big)\cdot\big(\nabla_{\theta}r_{w}-\nabla_{\theta}r_{l}\big),$ (17)
@@
    $=\Bigl(\underbrace{1-\sigma(d)}_{\text{BT magnitude}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha}\Bigr)\,\bigl\|\,\nabla_{\theta}r_{w}-\nabla_{\theta}r_{l}\,\bigr\|.$ (18)
@@
B.2 Training Details.
We implement all methods based on transformers (Wolf et al., 2020) and trl (von Werra et al., 2020), developing from the work of GRM (Yang et al., 2024). To use the Unified-Feedback dataset, we downsample the training data from the “all” set. All reward models are trained with the default reward head, as a linear layer with shape (hidden size, 1). For label smoothing, $\alpha$ is set to 0.1 by default. For all BT baselines, we conduct an extensive grid search over learning rates and report the best-tuned results. For BT-based variants (margin, margin-out, label smoothing), we search over a range of learning rates centered around the best value found for BT and report the best configuration. All models are trained for 1 epoch with full parameter tuning. We truncate the inputs over 4096 tokens, and an effective batch size of 256 with gradient accumulation. Our experiments are conducted using NVIDIA RTX A6000 49G.
@@
Appendix C Experiment Error Analysis
To assess the robustness of our main results reported in Table 1, we conduct an expanded evaluation on five random seeds for each method. Using a 40K subset of the Unified-Feedback dataset, we train both NormBT and the BT baseline following the same training configuration described in Appendix B. We compare these 10 models to show the performance improvement of NormBT over the BT baseline. The results reported in Table 4 and Figure 5 are consistent with the main experiments, with higher average scores across RewardBench and particularly strong gains in the Reasoning category.
Table 4: Error Analysis on Ten Models. Performance on RewardBench over models trained on five random seeds.
Reward Model Chat Chat Hard Safety Reasoning Average
BT (Baseline) 94.69 37.48 75.11 65.34 68.15
NormBT (Ours) 95.03 38.00 74.95 70.50 69.62
Improvement +0.34% ($\pm$0.19) +0.52% ($\pm$0.40) -0.16% ($\pm$0.49) +5.16% ($\pm$2.13) +1.46% ($\pm$0.61)
@@ Appendix C Experiment Error Analysis
Figure 5: Error Analysis on Ten Models. Detailed performance of models trained on five random seeds with standard errors on (a) RewardBench categories, and (b) Unified-Feedback Evaluation.
Appendix D Additional Examples
@@
Pair 1 (Go) Prompt. Write a Go function func LargestPrimeFactor(n int) int to solve the following problem: Return the largest prime factor of n. Assume n >> 1 and is not a prime.
>>> LargestPrimeFactor(13195)
29
>>> LargestPrimeFactor(2048)
2
isPrime := func(n int) bool {
    for i := 2; i < int(math.Pow(float64(n), 0.5)+1); i++ {
        if n%i == 0 {
            return false
    return true
largest := 1
for j := 2; j < n + 1; j++ {
    if n % j == 0 && isPrime(j) {
        if j > largest {
            largest = j
return largest
isPrime := func(n int) bool {
    for i := 2; i < int(math.Pow(float64(n), 0.5)+1); i++ {
        if n%i == 0 {
            return false
    return true
largest := 1
for j := 2; j < n + 1; j++ {
    if n % j == 0 && isPrime(n) {
        if j > largest {
            largest = j
return largest
Pair 2 (JavaScript) Prompt. Write a JavaScript function const isBored = (S) to solve the following problem. You’$l$ be given a string of words, and your task is to count the number of boredoms. A boredom is a sentence that starts with the word "I". Sentences are delimited by ’.’, ’?’ or ’!’.
let t = 0
if (S[0] == ’I’ && S[1] == ’ ’) { t = 1 }
for (let i = 0; i < S.length; i++) {
    if (S[i] == ’.’ $|$ S[i] == ’!’ $|$ S[i] == ’?’) {
        if (S[i + 1] == ’ ’ && S[i + 2] == ’I’ && S[i + 3] == ’ ’) {
            t++
return t
let t = 0
if (S[0] == ’ ’ && S[1] == ’I’) { t = 1 }
for (let i = 0; i < S.length; i++) {
    if (S[i] == ’.’ $|$ S[i] == ’!’ $|$ S[i] == ’?’) {
        if (S[i + 1] == ’ ’ && S[i + 2] == ’I’ && S[i + 3] == ’ ’) {
            t++
return t
Pair 3 (Go) Prompt. Write a Go function func NumericalLetterGrade(grades []float64) []string to solve the following problem. It is the last week of the semester and the teacher has to give the grades to students. The teacher has been making her own algorithm for grading, but she has lost the code she used for grading. You are given a list of GPAs and must output a list of letter grades using the following table:
GPA   | Letter grade
4.0   | A+
0.0   | E
letter_grade := make([]string, 0, len(grades))
for _, gpa := range grades {
    switch {
    case gpa == 4.0:
        letter_grade = append(letter_grade, "A+")
    case gpa > 3.7:
        letter_grade = append(letter_grade, "A")
    case gpa > 3.3:
        letter_grade = append(letter_grade, "A-")
    case gpa > 3.0:
        letter_grade = append(letter_grade, "B+")
    case gpa > 2.7:
        letter_grade = append(letter_grade, "B")
    case gpa > 2.3:
        letter_grade = append(letter_grade, "B-")
    case gpa > 2.0:
        letter_grade = append(letter_grade, "C+")
    case gpa > 1.7:
        letter_grade = append(letter_grade, "C")
    case gpa > 1.3:
        letter_grade = append(letter_grade, "C-")
    case gpa > 1.0:
        letter_grade = append(letter_grade, "D+")
    case gpa > 0.7:
        letter_grade = append(letter_grade, "D")
    case gpa > 0.0:
        letter_grade = append(letter_grade, "D-")
    default:
        letter_grade = append(letter_grade, ”E”)
return letter_grade
letter_grade := make([]string, 0, len(grades))
for _, gpa := range grades {
    switch {
    case gpa == 4.0:
        letter_grade = append(letter_grade, "A+")
    case gpa > 3.7:
        letter_grade = append(letter_grade, "A")
    case gpa > 3.3:
        letter_grade = append(letter_grade, "A-")
    case gpa > 3.0:
        letter_grade = append(letter_grade, "B+")
    case gpa > 2.7:
        letter_grade = append(letter_grade, "B")
    case gpa > 2.3:
        letter_grade = append(letter_grade, "B-")
    case gpa > 2.0:
        letter_grade = append(letter_grade, "C+")
    case gpa > 1.7:
        letter_grade = append(letter_grade, "C")
    case gpa > 1.3:
        letter_grade = append(letter_grade, "C-")
    case gpa > 1.0:
        letter_grade = append(letter_grade, "D+")
    case gpa > 0.7:
        letter_grade = append(letter_grade, "D")
    case gpa > 0.0:
        letter_grade = append(letter_grade, "D-")
    default:
        letter_grade = append(letter_grade, ”E+”)
return letter_grade
Appendix E Intuition: Large- vs. Small-Distance Pairs
Small representation distance can arise from two reasons: (1) uninformative or noisy pairs, such as near-duplicate responses that convey little preference signal; or (2) difficult but meaningful pairs, where the backbone fails to separate two responses despite clear human preference signal (e.g., correct vs. subtly flawed code). As illustrated in Figure 2(b), such case (2) pairs are common in practice, and Section D provides concrete examples.
In our datasets (UnifiedFeedback and Skywork-80K), most pairs are human-annotated or carefully curated, and following (Yang et al., 2024), pairs with identical ratings are filtered out. This substantially reduces the prevalence of case (1), making small-distance pairs far more likely to reflect case (2) of the difficult, fine-grained distinctions that RM should learn. Viewed through this lens, NormBT effectively upweights hard, meaningful pairs whose gradients would otherwise be suppressed due to representation limitations in the current model.
Appendix F Full vs. Proxy Representation Distance
(c) gemma-2b-it
(d) Llama-3.2-3b-Insturct
(e) Llama (with training)